Problem 8-2A (Video)

Lovell Computer Parts Inc. is in the process of setting a selling price on a new component it has just designed and developed. The following cost estimates for this new component have been provided by the accounting department for a budgeted volume of 54,000 units.
Per Unit Total
Direct materials $55
Direct labor $29
Variable manufacturing overhead $25
Fixed manufacturing overhead $648,000
Variable selling and administrative expenses $13
Fixed selling and administrative expenses $378,000

Lovell Computer Parts management requests that the total cost per unit be used in cost-plus pricing its products. On this particular product, management also directs that the target price be set to provide a 27% return on investment (ROI) on invested assets of $1,000,000.
Compute the markup percentage and target selling price that will allow Lovell Computer Parts to earn its desired ROI of 27% on this new component. (Round markup percentage to 2 decimal places, e.g. 10.50%.)
Markup percentage

%
Target selling price $

Assuming that the volume is 43,200 units, compute the markup percentage and target selling price that will allow Lovell Computer Parts to earn its desired ROI of 27% on this new component. (Round answers to 2 decimal places, e.g. 10.50% or 10.50.)
Markup percentage

%
Target selling price $
Answer #1
1
Direct materials 55
Direct labor 29
Variable manufacturing overhead 25
Fixed manufacturing overhead 12 =648000/54000
Variable selling and administrative expenses 13
Fixed selling and administrative expenses 7 =378000/54000
Total cost per unit 141
Desired ROI per unit 5 =(1000000*27%)/54000
Desired ROI per unit 5
Divide by Total cost per unit 141
Markup Percentage 3.55%
Desired ROI per unit 5
Add: Total cost per unit 141
Target Selling price 146
2
Direct materials 55.00
Direct labor 29.00
Variable manufacturing overhead 25.00
Fixed manufacturing overhead 15.00 =648000/43200
Variable selling and administrative expenses 13.00
Fixed selling and administrative expenses 8.75 =378000/43200
Total cost per unit 145.75
Desired ROI per unit 6.25 =(1000000*27%)/43200
Desired ROI per unit 6.25
Divide by Total cost per unit 145.75
Markup Percentage 4.29%
Desired ROI per unit 6.25
Add: Total cost per unit 145.75
Target Selling price 152.00